# Growth of Light-Seed Black Holes in Gas-rich Galaxies at High Redshift

### v1

@@ 2.2 Isolated Galaxy Evolution
Once the initial conditions have been created and read in by Arepo , the galaxy is evolved initially for $100$ million years at a resolution of $1$ $\rm{pc}$ without sink particle routines turned on. This is done to ensure the galaxy has a suitably chaotic nature, similar to what is observed for high-z galaxies. Within this period, the spherical bulge of gas cells collapses to form a chaotic planar structure - similar to high-z galaxies. The state of the galaxies at this time can be seen in Figure 1. Once the simulations have been evolved for 100 Myrs, we then restart the simulation at our canonical resolution of $10^{-2}$ $\rm{pc}$ with sink particle routines turned on (see §2.3). We perform a total of two separate realisations - one with SNe feedback turned off and the other with SNe feedback turned on. The high resolution ($10^{-2}$ $\rm{pc}$) of our simulation prohibits us from running the simulations for extended periods and so we evolve the simulations for only 10 Myrs. However, this is justified as the free-fall time of our halo is approximately 5 Myrs and we also do not account for cosmological (major) mergers which would be expected to occur on a timescale of a few times the dynamical time of the halo.
Figure 1: Left hand panel: The parent galaxy that we use in both of our simulations (feedback and no-feedback). This is the state of the galaxy immediately before we evolve it at our highest resolution ($\Delta x=10^{-2}$ pc) Right hand panel: The same image but zoomed in to the densest gas clump of the galaxy.
@@ 2.3 Sink Particle Technique
The sink particle technique employed in this work builds and improves on the scheme originally introduced into Arepo by Wollenberg et al. (2020) and Tress et al. (2020). Sink particles are formed in high density gas regions when a gas cell can no longer refine upon reaching the minimum cell size of the mesh. They are point like particles artificially introduced to model the behaviour of physical processes that occur below the resolution of our mesh. The original scheme introduced into Arepo converts a cell directly into a sink particle once three conditions are met: 1) the cell reaches a user defined threshold density 2) the cell is sufficiently far away from a pre-existing sink so that their accretion radii do not overlap and 3) the gas within this region is gravitationally bound and collapsing. The approach taken here is broadly similar with the criteria extended somewhat based on the implementation of Krumholz et al. (2004) and previously introduced into the Enzo code (Bryan et al., 2014; Brummel-Smith et al., 2019) by Regan & Downes (2018). In this case, we change the formation criteria so that a cell is converted into a sink only after the Jeans density threshold is cleared. This removes the user-defined threshold and in doing so makes the sink formation more self-consistent. The new formation criteria checks for conditions 2, 3 as mentioned above along with these additional conditions:
@@
    We first check that the candidate cell is at the highest refinement level and can no longer refine.
@@
    where $\Delta x$ is the size of the cell and $J=0.25$ is a constant (Truelove et al., 1997).
This technique, as noted above follows the formation technique introduced in Regan & Downes (2018) (dubbed SmartStars). In addition to the augmented formation criteria, the type of sink particle formed can be discriminated into four different flavours namely: population II (PopII), population III (PopIII), super-massive star (SMS), and massive black hole (MBH). For the purposes of this study, we only allow (metal-free) PopIII star formation followed by the transition of the sink into a MBH particle where appropriate (see §2.5). Henceforth, whenever a sink particle is in its stellar phase, we call it a PopIII particles and after the stellar lifetime has ended, we call it a MBH particle.
@@ 2.4 Assigning Sink Particle Masses
Upon the formation of a PopIII particle, the mass of the corresponding host cell is directly transferred to the newly created PopIII particle, effectively converting the host cell into the particle itself, as outlined in the preceding discussion. This is the PopIII particle’s dynamical mass. However, we need to be careful here. Our spatial resolution ($\Delta x=10^{-2}$ $\rm{pc}$), while good, is far from the scale of actual star formation. As a result, we additionally assign stellar masses to each PopIII particle. The stellar masses are randomly sampled from a top-heavy initial mass function (IMF).
@@
To construct the IMF, we set the upper and lower limit at $300\ \text{M}_{\odot}$ and $1\ \text{M}_{\odot}$, respectively and its slope is $-1.3$. We set the characteristic mass, $\rm{M_{char}}$, to be 20 $\rm{M_{\odot}}$. Each PopIII particle is also assigned a stellar life according to its stellar mass. The stellar lifetimes of PopIII stars ranges from a few times $10^{6}$ to a few times $10^{7}$ years, which is of the order of our entire simulation. To properly capture mass growth through accretion - which is the primary goal of this study - we divide the stellar lifetimes by $100$. This modification also allows us to model the fact that some stars would be brought into the galaxy via mergers - a process not accounted for in our modelling.
At the end of its lifetime, each PopIII particle must transition to either a MBH particle or be deleted following a Pair Instability SNe (PISN) event depending on whether or not feedback is turned on. If feedback is turned on, the PopIII particle may experience a SNe (depending on its stellar mass) and inject energy into the surrounding (see §2.5 for more info) gas. If a PopIII particle is in the stellar mass range where it should transition to a MBH then the MBH can accrete the surrounding gas following its formation. Only MBH particles can accrete. Mergers between MBHs are also possible via our framework but in this study we choose not to implement MBH mergers and instead treat them as collisionless particles. We leave a study of MBH mergers in high-z galaxies to a follow-up work.
@@ 2.5 Supernova Feedback
After the assigned stellar lifetime of a PopIII particle has passed, it may undergo a SNe explosion depending on its stellar mass. Following a SNe explosion, the PopIII particle either transitions into a MBH particle or is deleted (following a PISN event). The SNe categories and energy releases are as follows:
@@
    PopIII particles in the mass range of 20-40 $\ \text{M}_{\odot}$ undergo a hypernova. We extrapolate values of $E_{\rm{SN}}$ and mHe from Nomoto et al. 2006. PopIII particles in this range transition into a MBH following the explosion. The MBH mass is set to the helium core mass post explosion.
@@
    PopIII particles in the mass range of $40-140$ $\ \text{M}_{\odot}$ experience direct collapse into a BH and exhibit no SNe explosion. The total mass of the PopIII is converted into the MBH particle.
@@
Unlike previous works that ensure the Sedov-Taylor scale is resolved so that the energy of the SNe can be directly injected into the simulation mesh (see e.g. Gatto et al. 2015; Tress et al. 2020; Magg et al. 2022), we instead opt to model the explosion as an injection of momentum. As in Magg et al. (2022), we inject the momentum into the nearest 1000 cells surrounding the PopIII particle. At our resolution, this corresponds to an injection region of $\sim$ 0.1 pc. We calculate the momentum of the blast according to the energy of the SNe, density and metallicity of the injection region as (Smith et al., 2019).
@@ 2.6 Sink Particle Accretion
Once a PopIII particle has turned into a MBH particle, we accrete onto it using the Bondi-Hoyle-Lyttleton (BHL) (Bondi, 1952) method following the prescription given in Krumholz et al. (2004). Accretion onto MBH particle then occurs by removing gas from the surrounding accretion sphere. The size of the accretion sphere is 4 cells wide.
@@
where $\rho_{\infty}$ is the weighted density inside the accretion sphere, $r_{BH}=\frac{GM_{sink}}{v_{\infty}^{2}+c_{\infty}^{2}}$ is the BHL radius, $v_{\infty}$ is the mass weighted velocity and $c_{\infty}$ is the sound speed of the surrounding gas. For $v_{\infty}$, we take the relative velocity of the MBH particle and the host cell and for $c_{\infty}$, we take the local sound speed. For calculating the value of $\rho_{\infty}$, we need to be more careful and here we follow the procedure outlined in Krumholz et al. (2004). The value of $\rho_{\infty}$ is calculated by assigning weights to all cells within the kernel radius, $r_{K}$. The kernel radius is given by:
@@
For all but the very smallest black hole masses (100 $\rm{M_{\odot}}~{}$ $\sim 0.1$ pc BHL radius) we can resolve the BHL radius with our resolution. Once the kernel radius is determined, we use this to assign a weight to every cell within our accretion sphere using the formula:
@@
where $\bar{\rho}$ is the mass weighted mean density within the accretion sphere. It is this value of $\rho_{\infty}$ that is then used in calculating the accretion rate (Eqn 7) at each timestep. In order to improve our scheme further, we also include the impact of vorticity on our accretion rate.
@@ 2.7 Vorticity
Given the somewhat chaotic nature of high-z galaxies, we also take into account the vorticity, $\omega$, of the surrounding gas. To determine the vorticity of the gas, we calculate the mass-weighted curl of all the gas cells following the prescription set out in Krumholz et al. (2006). The vorticity, $\omega$, is defined as
@@
thewith the dimensionless vorticity $\omega_{*}$ given by
@@
For a given timestep, $t_{h}$, the MBH particle’s mass is increased by $t_{h}*\dot{M}$. For each gas cell, we calculate the fraction of mass lost to the MBH particle. The accretion scheme is constructed in such a way as to explicitly conserve linear momentum between the gas cells and the MBH particle.
@@ 2.8 Chemistry solver
We model the chemistry using SGCHEM (Hartwig et al., 2015) as recently employed by Wollenberg et al. (2020) and Prole et al. (2022), which consists of 12 species, H, H+, H-, H2+ , H2 , He, He+, He++, D, D+, HD, and e-. The primordial gas follows 45 reactions based on the chemical network described in Clark et al. (2011) and updated by Schauer et al. (2017). The initial gas fractions are $2\times 10^{-6}$ for H2, $10^{-4}$ for H+, $2.6\times 10^{-9}$ for D+, and $2.6\times 10^{-5}$ for D (the fraction is with respect to neutral hydrogen). The chemical network accounts for H2 cooling, HD cooling, heating and cooling from gas, shocks, compression, and expansion of the gas, collisionally induced H2 emission, ionization and recombination, and also heating from accretion luminosity. The adiabatic index of the gas is computed as a function of chemical composition and temperature using the Arepo Harten-Lax-van Leer with contact and Alfvén HLLD Riemann solver.
@@ 3 Results
Figure 2: Sink formation (black dots) in a box of $5$ pc centered at the densest clump of the galaxy at T = $40$ kyr. Even at this early stage in the evolution multiple PopIII particles have formed in the densest regions.
In Fig. 1, we show the state of the galaxy after allowing the galaxy to evolve at lower resolution for 100 Myr after the initial setup. We set the time at this point to be T = 0. From this point onwards, the simulation box is evolved at a resolution of $\Delta x=10^{-2}$ pc. The galaxy clearly shows a chaotic structure typical of high-z galaxies. The top-right panel is a zoom-in image at the same time around the densest gas cell in the galaxy. The gas cloud shown here, along with multiple other clouds, form spiral arms due to the rotation of the galaxy. This served as the starting point for the next phase of the simulations, which we henceforth refer to in the cases where SNe feedback was and was not enabled as the ‘feedback’ and ‘no-feedback’ simulations, respectively. Note that all PopIII particles are converted into MBH particles following their stellar lifetimes in the no-feedback case, while their conversion or deletion (in the case of PISN) depends on their stellar mass in the feedback simulation.
In Fig. 2, we show a zoom-in image at $t\sim 40$ kyr for the no-feedback simulation. Annotated on top of the image are the location of multiple sink particles - in this case PopIII particles. In the image, we see the formation of several PopIII particles at the center of fragmenting gas clouds. Recall that a PopIII particle has the same dynamical mass as its progenitor gas cell but is assigned a stellar mass by randomly sampling an IMF. These PopIII particles continue to evolve with more PopIII particles forming over the course of the simulation. As this is for the no-feedback simulation, no SNe explosions have occurred or will occur.
@@
In Fig. 3, the initial PopIII particle masses are shown from both the no-feedback and feedback simulations. We form $2004$ and $2181$ PopIII particles in the no-feedback and feedback simulations, respectively, over the course of 10 Myrs. A bimodal distribution is observed extending from $\sim 30$ to $\sim 10^{3}$ $\ \text{M}_{\odot}$ with the two peaks around $\sim 200$ $\rm{M_{\odot}}~{}$and $\sim 400$ $\ \text{M}_{\odot}$. The second peak PopIII particles are created from higher temperature gas cells, which has a higher Jeans density thereby impacting on the sink formation criteria via the Jeans density threshold. This means that higher temperature gas can lead to higher sink masses as expected. Consequently, in the feedback simulation, we get a higher number of massive PopIII particles and the bi-modal distribution is more prominent. In both cases, these are the dynamical masses of the PopIII particles. We now examine the results from the no-feedback and feedback simulations in turn.
3.1 No-feedback Simulation
@@
Once a PopIII particle reaches the end of its stellar lifetime, it is converted into a MBH particle and it is able to accrete gas. We show the growth of MBH particles through accretion in Fig. 4 for the no-feedback case. In the left panel we only show MBH particles that accrete more than $1\ \text{M}_{\odot}$ over the course of the simulation - we henceforth refer to these as active sinks. The stellar phase of PopIII particles is not shown in this case, i.e., $t=0$ refers to the time at which the PopIII particle was converted into a MBH particle. Initially, accretion is suppressed as excess angular momentum must be shed before accretion can accelerate. However, as accretion occurs, the angular momentum of the gas slowly decreases, which increases the accretion rate, which in turn decreases angular momentum. As a result, a rapid run-away accretion episode occurs, which increases the gas mass to an excess of $10^{5}$ $\ \text{M}_{\odot}$ for the most rapidly accreting MBH particles. A total of $1520$ MBH particles out of $2004$ accrete more than 1 solar mass (i.e. approximately 75%). Of this total number of actively accreting seeds, 33 reach final masses in excess of $10^{5}\ \text{M}_{\odot}$. The most-massive MBH particle reaches a final mass of $\sim 7.5\times 10^{5}$ $\ \text{M}_{\odot}$.
In the right hand panel, we show the Eddington ratio for the most rapidly accreting MBH particles - those that reach masses in excess of $10^{5}$ $\rm{M_{\odot}}~{}$in less than 10 Myrs. The accretion is initially in the hyper-Eddington regime but as the surrounding gas mass is gradually depleted the accretion rate falls to values closer to the Eddington rate as the initially plentiful supply of gas is consumed.
Figure 5: Final mass of the MBH particles v/s total gas mass within $r_{BH}$ of the MBH particles. The total gas mass is calculated from the entire simulation of $10^{7}$ years. The color bar shows the average density of the gas. The MBH particles that undergo runaway accretion capture $\sim 1\%$ of the total gas mass.

### v2

@@ 2.2 Isolated Galaxy Evolution
Once the initial conditions have been created and read in by Arepo , the galaxy is evolved initially for $100$ million yr at a resolution of $1$ $\rm{pc}$ without the sink particle routines turned on. This was done to ensure that the galaxy has a suitably chaotic nature, similar to what is observed for high-z galaxies. Within this period, the spherical bulge of gas collapses to form a turbulent planar structure. The state of the galaxies at this time can be seen in Figure 1. Once the simulations have evolved for 100 Myr, we then restart the simulation at our canonical resolution of $10^{-2}$ $\rm{pc}$ cell size with the sink particle routines turned on (see §2.3). We perform a total of two separate realisations- one with SNe feedback turned off and the other with SNe feedback turned on. The high resolution ($10^{-2}$ $\rm{pc}$) of our simulation prohibits us from running the simulations for extended periods, so we evolve the simulations for only 10 Myr. However, this is justified as the free-fall time of our halo is approximately 5 Myr, and we also do not account for cosmological (major) mergers, which would be expected to occur on a timescale of a few times the dynamical time of the halo.
@@ 2.3 Sink Particle Technique
The sink particle technique employed in this work builds and improves on the scheme originally introduced into Arepo by Wollenberg et al. (2020) and Tress et al. (2020). Sink particles are formed in high density gas regions when a gas cell can no longer refine upon reaching the minimum cell size of the mesh. They are point like particles artificially introduced to model the behaviour of physical processes that occur below the resolution of our mesh. The original scheme introduced into Arepo converts a cell directly into a sink particle once three conditions are met: 1) the cell reaches a user defined threshold density 2) the cell is sufficiently far away from a pre-existing sink so that their accretion radii do not overlap, and 3) the gas within this region is gravitationally bound and collapsing. The approach taken here is broadly similar with the criteria extended somewhat based on the implementation of Krumholz et al. (2004) and previously introduced into the Enzo code (Bryan et al., 2014; Brummel-Smith et al., 2019) by Regan & Downes (2018). In this case, we change the formation criteria so that a cell is converted into a sink only after the Jeans density threshold is cleared. This removes the user-defined threshold and in doing so makes the sink formation more self-consistent. The new formation criteria checks for conditions 2, 3 as mentioned above along with these additional conditions:
@@
    We first check that the candidate cell is at the highest refinement level and can no longer be refined.
@@
    where $\Delta x$ is the cell size and $J=0.25$ is a constant (Truelove et al., 1997).
This technique, as noted above, follows the formation technique introduced in Regan & Downes (2018) (dubbed SmartStars). In addition to the augmented formation criteria, the type of sink particle formed can be discriminated into four different flavours, namely: population II (PopII), population III (PopIII), super-massive star (SMS), and massive black hole (MBH). For the purposes of this study, we only allow (metal-free) PopIII star formation followed by the transition of the sink into a MBH particle where appropriate (see §2.5). Henceforth, whenever a sink particle is in its stellar phase, we call it a PopIII particle, and after the stellar lifetime has ended, we call it a MBH particle.
@@ 2.4 Assigning Sink Particle Masses
Upon the formation of a PopIII particle, the mass of the corresponding host cell is directly transferred to the newly created PopIII particle, effectively converting the host cell into the particle itself, as outlined in the preceding discussion. This is the dynamical mass of the PopIII particle. However, we need to be careful here. Our spatial resolution ($\Delta x=10^{-2}$ $\rm{pc}$), while good, is far from the scale of actual star formation. As a result, we additionally assign stellar masses to each PopIII particle. The stellar masses are randomly sampled from a top-heavy initial mass function (IMF).
@@
To construct the IMF, we set the upper and lower limits at $300\ \text{M}_{\odot}$ and $1\ \text{M}_{\odot}$, respectively, and its slope is $-1.3$. We set the characteristic mass, $\rm{M_{char}}$, to 20 $\rm{M_{\odot}}$. We choose a lower limit of $1\ \text{M}_{\odot}$ due to the constraint of cell size resolution. At a maximum resolution of $10^{-2}$ pc, we are unable to resolve structures significantly below $1\ \text{M}_{\odot}$ and hence we set this as the lower limit for star formation in cells. Each PopIII particle is also assigned a stellar life according to its stellar mass. The stellar lifetimes of PopIII stars range from a few times $10^{6}$ to a few times $10^{7}$ yr, which is of the order of our entire simulation. To account for this and allow BH remnants to form sufficiently rapidly, we divide the stellar lifetimes by 100. This modification can also be thought of to account for the previous generation of star formation missed through our idealised setup. We also do not account for radiative feedback from the PopIII stars in this work (although we do account for supernova feedback). This is a limitation of our work and is discussed in detail in §5
At the end of its lifetime, each PopIII particle must transition to either a MBH particle or be deleted following a Pair Instability SNe (PISN) event depending on whether or not feedback is turned on. In the case of no-feedback simulation, all PopIII particles are converted to MBH particles including those that would potentially undergo PISN. If feedback is turned on, the PopIII particle may experience a SNe (depending on its stellar mass) and inject energy into the surrounding gas (see §2.5 for more information). If a PopIII particle is in the stellar mass range where it should transition to a MBH then the MBH can accrete the surrounding gas following its formation. Only MBH particles can accrete. Mergers between MBHs are also possible via our framework, but in this study we choose not to implement MBH mergers and instead treat them as collisionless particles. We leave a study of MBH mergers in high-z galaxies to a follow-up work.
Figure 1: Left hand panel: Face-on projection of the parent galaxy that we use in both of our simulations (feedback and no-feedback). The projection is created using the density weighing scheme. This is the state of the galaxy immediately before we evolve it at our highest resolution ($\Delta x=10^{-2}$ pc). Sink formation is not activated at this point. Right hand panel: The same image but zoomed in to the densest gas clump of the galaxy.
Figure 2: Left hand panel: Edge-on projection of the parent galaxy that we use in both of our simulations (feedback and no-feedback). This visualisation is identical to Figure 1 apart from the orientation.
@@ 2.5 Supernova Feedback
After the stellar lifetime assigned to a PopIII particle has passed, it may undergo a SNe explosion depending on its stellar mass. Following a SNe explosion, the PopIII particle either transitions into a MBH particle or is deleted (following a PISN event). The SNe categories and energy releases are as follows:
@@
    PopIII particles in the mass range of 20-40 $\ \text{M}_{\odot}$ undergo a hypernova. We extrapolate the values of $E_{\rm{SN}}$ and mHe from Nomoto et al. 2006. PopIII particles in this range transition to a MBH following the explosion. The MBH mass is set to the helium core mass post explosion.
@@
    PopIII particles in the mass range of $40-140$ $\ \text{M}_{\odot}$ experience a direct collapse into a BH and exhibit no SNe explosion. The total mass of the PopIII particle is converted into the MBH particle.
@@
Unlike previous works that ensure the Sedov-Taylor scale is resolved so that the energy of the SNe can be directly injected into the simulation mesh (see, e.g. Gatto et al. 2015; Tress et al. 2020; Magg et al. 2022), we instead opt to model the explosion as an injection of momentum. As in Magg et al. (2022), we inject the momentum into the nearest 1000 cells surrounding the PopIII particle. At our resolution, this corresponds to an injection region of $\sim$ 0.1 pc. We calculate the momentum of the blast according to the SNe energy, the density and the metallicity of the injection region as (Smith et al., 2019).
@@ 2.6 Sink Particle Accretion
Once a PopIII particle has turned into a MBH particle, we accrete onto it using the Bondi-Hoyle-Lyttleton (BHL) (Bondi, 1952) method following the prescription given in Krumholz et al. (2004). Accretion onto a MBH particle then occurs by removing gas from the surrounding accretion sphere. The size of the accretion sphere is 4 cells wide.
@@
where $\rho_{\infty}$ is the weighted density inside the accretion sphere, $r_{BH}=\frac{GM_{sink}}{v_{\infty}^{2}+c_{\infty}^{2}}$ is the BHL radius, $v_{\infty}$ is the mass weighted velocity and $c_{\infty}$ is the sound speed of the surrounding gas. For $v_{\infty}$, we take the relative velocity of the MBH particle and the host cell, and for $c_{\infty}$, we take the local sound speed. To calculate the value of $\rho_{\infty}$, we need to be more careful, and here we follow the procedure outlined in Krumholz et al. (2004). The value of $\rho_{\infty}$ is calculated by assigning weights to all cells within the kernel radius, $r_{K}$. The kernel radius is given by:
@@
For all but the very smallest black hole masses (100 $\rm{M_{\odot}}~{}$ $\sim 0.1$ pc BHL radius), we can resolve the BHL radius with our resolution. Once the kernel radius is determined, we use this to assign a weight to every cell within our accretion sphere using the formula:
@@
where $\bar{\rho}$ is the mass weighted mean density within the accretion sphere. This value of $\rho_{\infty}$ is used in calculating the accretion rate (Eqn 7) at each timestep. In order to improve our scheme further, we also include the impact of vorticity on our accretion rate.
@@ 2.7 Vorticity
Given the somewhat chaotic nature of high-z galaxies, we also take into account the vorticity, $\omega$, of the surrounding gas. To determine the vorticity of the gas, we calculate the mass-weighted curl of all gas cells following the prescription set out in Krumholz et al. (2006). The vorticity, $\omega$, is defined as
@@
with the dimensionless vorticity $\omega_{*}$ given by
@@
For a given time step, $t_{h}$, the mass of the MBH particle increases by $t_{h}*\dot{M}$. For each gas cell, we calculate the fraction of mass lost to the MBH particle. The accretion scheme is constructed in such a way as to explicitly conserve the linear momentum between the gas cells and the MBH particle.
As a BH accretes, it will also release kinetic, thermal, and radiative energy in the form of jets or winds - so-called BH feedback. In this study, we have ignored this feedback mechanism and this caveat is discussed in detail in §5. We will explore the impact of BH feedback on our results in a forthcoming study.
@@ 2.8 Chemistry solver
We model the chemistry using SGCHEM (Hartwig et al., 2015) as recently employed by Wollenberg et al. (2020) and Prole et al. (2022), which consists of 12 species, H, H+, H-, H2+ , H2 , He, He+, He++, D, D+, HD, and e-. The primordial gas follows 45 reactions based on the chemical network described in Clark et al. (2011) and updated by Schauer et al. (2017). The initial gas fractions are $2\times 10^{-6}$ for H2, $10^{-4}$ for H+, $2.6\times 10^{-9}$ for D+, and $2.6\times 10^{-5}$ for D (the fraction is with respect to neutral hydrogen). The chemical network accounts for H2 cooling, HD cooling, heating and cooling from gas, shocks, compression, and expansion of the gas, collisionally induced H2 emission, ionisation, and recombination, and also heating from accretion luminosity. The adiabatic index of the gas is computed as a function of the chemical composition and temperature using the Arepo Harten-Lax-van Leer with contact and Alfvén HLLD Riemann solver.
@@ 3 Results
Figure 3: Star formation efficiency (SFE) for both the no-feedback (blue) and feedback (orange) simulations. We see a rise in SFE at the beginning, converting almost 0.1% of the gas to stars. The SFE for feedback simulation is higher in the first $10^{6}$ yr overall. Nonetheless, the galaxy remains gas rich.
Figure 4: Sink formation (black dots) in a box of $5$ pc centred at the densest clump of the galaxy at T = $40$ kyr. The black dot represents PopIII particles. Even at this early stage in the evolution multiple PopIII stars have formed in the densest regions.
In Figs. 1 (face-on) and 2 (edge-on), we show the state of the galaxy after allowing the galaxy to evolve at lower resolution for 100 Myr after the initial setup. We set the time at this point to be T = 0. From this point on, the simulation box is evolved at a resolution of $\Delta x=10^{-2}$ pc. The galaxy clearly shows a turbulent planar structure. The top-right panel in both figures is a zoom-in image at the same time around the densest gas cell in the galaxy. The gas cloud shown here, along with multiple other clouds, forms spiral arms due to the rotation of the galaxy. This serves as the starting point for the next phase of the simulations, which we henceforth refer to in the cases where SNe feedback was and was not enabled as the ‘feedback’ and ‘no-feedback’ simulations, respectively. Note that all PopIII particles are converted into MBH particles following their stellar lifetimes in the no-feedback case, while their conversion or deletion (in the case of PISN) depends on their stellar mass in the feedback simulation. We show the star formation efficiency (SFE), defined as the mass converted to PopIII particles per unit gas mass, in Fig. 3 for both simulations. We see a sharp increase in SFE in the first $10^{5}$ yr reaching higher than 0.1% for both simulations. Our ICs have multiple dense gas clumps, ideal for star formation bursts, resulting in this sharp increase. Once the available dense gas converts into PopIII particles, the SFE decreases. We also observe a small bump in SFE between 2-3×$\times 10^{5}$ yr. At this point, the majority of PopIII particles have been converted to MBH particles, which accrete a lot of gas. The large scale gas movement bumps the SFE to 0.02-0.04 %. Around the same time in the feedback simulation, a lot of PopIII particles undergo supernova. This feedback pushes and disturbs the gas, creating conditions for further star formation episodes. We see that the SFE increases to 0.06-0.07% in the feedback simulation as a result. After $4\times 10^{5}$ yr in both simulations, the SFE becomes smaller than $10^{-3}\%$. Overall, our SFEs remain fairly low (much less than 1%).
In Fig. 4, we show a zoom-in image at $t\sim 40$ kyr for the no-feedback simulation. Annotated on top of the image are the locations of multiple sink particles - in this case PopIII particles. In the image, we see the formation of several PopIII particles at the centre of fragmenting gas clouds. Recall that a PopIII particle has the same dynamical mass as its progenitor gas cell but is assigned a stellar mass by randomly sampling an IMF. These PopIII particles continue to evolve, with more PopIII particles forming over the course of the simulation. As this is for the no-feedback simulation, no SNe explosions have occurred or will occur.
@@
In Fig. 5, the initial PopIII particle masses are shown from both the no-feedback and feedback simulations. We form $2004$ and $2181$ PopIII particles in the no-feedback and feedback simulations, respectively, over the course of 10 Myr. A bimodal distribution is observed extending from $\sim 30$ to $\sim 10^{3}$ $\ \text{M}_{\odot}$ with two peaks around $\sim 200$ $\rm{M_{\odot}}~{}$and $\sim 400$ $\ \text{M}_{\odot}$. The second peak PopIII particles are created from higher temperature gas cells, which has a higher Jeans density thereby impacting on the sink formation criteria via the Jeans density threshold. This means that higher temperature gas can lead to higher sink masses as expected. Consequently, in the feedback simulation, we get a higher number of massive PopIII particles and the bi-modal distribution is more prominent. In both cases, these are the dynamical masses of the PopIII particles. We now examine the results from the no-feedback and feedback simulations in turn.
3.1 No-feedback Simulation
Once a PopIII particle reaches the end of its stellar lifetime, it is converted into a MBH particle, and it is able to accrete gas. We show the growth of MBH particles through accretion in Fig. 6 for the no-feedback case. In the left panel we only show MBH particles that accrete more than $1\ \text{M}_{\odot}$ over the course of the simulation - we henceforth refer to these as active sinks. The stellar phase of PopIII particles is not shown in this case, i.e. $t=0$ refers to the time at which the PopIII particle was converted into a MBH particle. Initially, accretion is suppressed as excess angular momentum must be shed before accretion can accelerate. However, as accretion occurs, the angular momentum of the gas slowly decreases, which increases the accretion rate, which in turn decreases the angular momentum. As a result, a rapid run-away accretion episode occurs, increasing the gas mass to an excess of $10^{5}$ $\ \text{M}_{\odot}$ for the most rapidly accreting MBH particles. A total of $1520$ MBH particles out of $2004$ accrete more than 1 solar mass (i.e. approximately 75%). Of this total number of actively accreting seeds, 33 reach final masses in excess of $10^{5}\ \text{M}_{\odot}$. The most-massive MBH particle reaches a final mass of $\sim 7.5\times 10^{5}$ $\ \text{M}_{\odot}$.
In the right hand panel, we show the Eddington ratio for the most rapidly accreting MBH particles - those that reach masses in excess of $10^{5}$ $\rm{M_{\odot}}~{}$in less than 10 Myr. The accretion is initially in the hyper-Eddington regime, but as the surrounding gas mass is gradually depleted, the accretion rate falls to values closer to the Eddington rate as the initially plentiful supply of gas is consumed.
Figure 7: Final mass of the MBH particles v/s total gas mass within $r_{BH}$ of the MBH particles. The total gas mass is calculated from the entire simulation of $10^{7}$ yr. The colour bar shows the average density of the gas. The MBH particles that undergo runaway accretion capture $\sim 1\%$ of the total gas mass.
Figure 8: Time series projection plots for the supernova feedback simulation. All panels are centred at the densest clump in the IC file. The top panels show the entire galaxy and we zoom in to the densest clump in the bottom plots. The density projection is calculated the same way as Fig. 1. The black dots represent the PopIII and MBH particles. We see on global scales, supernova feedback causes shells of gas to expand radially. On smaller scales, the clumps are completely disrupted.